Please help I need to know if 34=-2(23-y)+4(y-4) has no solution, has all real numbers in the solution or the solution set is {?}
Answer by Boreal(15235)

34=-2(23-y)+4(y-4)
34= -46+2y+4y-16
34= -62+ 6y
96=6y
y=16 That is the solution.
check
34= -2(7) + 4 (12)
34=34